OpenStudy (anonymous):

I have read my math book and all and still don't understand this problem. 2/5 + ((5 − (5/9))/5/6 · 3/16

OpenStudy (campbell_st):

ok... start with the brackets whats 5 - 5/9 =

OpenStudy (anonymous):

40/9

OpenStudy (campbell_st):

ok... next the division 40/9 divided by 5/6 you need to know the rule for dividing by a fraction.

OpenStudy (anonymous):

thats where i get lost

OpenStudy (campbell_st):

ok... went dividing by a fraction, flip the fraction and multiply \[\frac{\frac{40}{9}}{\frac{5}{6}} = \frac{40}{9} \times \frac{6}{5}\] you'll need to evalute this

OpenStudy (anonymous):

do i need common denominators?

OpenStudy (campbell_st):

nope... you can cancel common factors... to simplify the fraction

OpenStudy (campbell_st):